Find the average weight in pounds of a cardinal. Use the rational exponent equation L =2.43* W^0.3326 to estimate the wingspan L in feet of the bird that weighs W pounds Include in your post the type of bird and the average weight and show the calculations necessary to find the approximate wingspan. Average weight is 2.96 ounces.
L = 2.43 * w^0.3326
L=2.43x .185^0.3326
L = 2.43
L =
I don’t understand where to go from here. 2.43 * 2.96 ^0.3326
Now what do I do with the ^0.3326?
2.96 = 32 ounces = .185

Answer by stanbon(75887)

Use the rational exponent equation L =2.43* W^0.3326 to estimate the wingspan L in feet of the bird that weighs W pounds Include in your post the type of bird and the average weight and show the calculations necessary to find the approximate wingspan. Average weight is 2.96 ounces.
---------------------
L = 2.43*2.96^(0.3326)
You need a calculator for this nex step.
L = 2.43*1.434668699...
Again, use a calculator to get:
L = 3.486244938...ft.